# Editing Knowledge Representation of Language Lodel via Rephrased Prefix Prompts

### v1

@@ Abstract
Neural language models (LMs) have been extensively trained on vast corpora to store factual knowledge about various aspects of the world described in texts. Current technologies typically employ knowledge editing methods or specific prompts to modify LM outputs. However, existing knowledge editing methods are costly and inefficient, struggling to produce appropriate text. Additionally, prompt engineering is opaque and requires significant effort to find suitable prompts. To address these issues, we introduce a new method called PSPEM (Prefix Soft-Prompt Editing Method), that can be used for a lifetime with just one training. It resolves the inefficiencies and generalizability issues in knowledge editing methods and overcomes the opacity of prompt engineering by automatically seeking optimal soft prompts. Specifically, PSPEM utilizes a prompt encoder and an encoding converter to refine key information in prompts and uses prompt alignment techniques to guide model generation, ensuring text consistency and adherence to the intended structure and content, thereby maintaining an optimal balance between efficiency and accuracy. We have validated the effectiveness of PSPEM through knowledge editing and attribute inserting. On the COUNTERFACT dataset, PSPEM achieved nearly 100% editing accuracy and demonstrated the highest level of fluency. We further analyzed the similarities between PSPEM and original prompts and their impact on the model’s internals. The results indicate that PSPEM can serve as an alternative to original prompts, supporting the model in effective editing.
@@ 1 Introduction
Figure 1: By inputting the prompts on the left side into the model, traditional prompt engineering generates text with erroneous information, while PSPEM can correct such errors.
Language models based on the Transformer architecture, such as GPT and BERT, have revolutionized various natural language processing tasks with their capacity to store and utilize real-world factual knowledge within their parameters [27, 8, 24]. For example, when asked, “Where is the Eiffel Tower located?” GPT provides the accurate answer, “Paris”. However, inconsistencies or biases present in the pre-training data can propagate into the text generated by the model, leading to errors or contradictions. Additionally, the knowledge of the world changes over time, which presents a challenge to the static knowledge in the model. Addressing this issue requires a nuanced approach to updating the model’s knowledge base. Merely retraining the entire model with new data is prohibitively expensive and time-consuming, while fine-tuning on specific updated knowledge risks overfitting and losing the model’s generalization capabilities.
@@
Methods of model editing fall into two distinct classifications depending on the alteration of the original model weights: weight-preserved and weight-modified methods [30]. Weight-preserved strategies typically necessitate the inclusion of extra content, while weight-modified techniques directly alter the model’s weights. Weight modification methods include hypernetwork-based learning methods and direct optimization methods. Hypernetwork-based learning methods, such as KE [7], MEND [21] and MALMEN [26], utilize a hypernetwork to predict essential updates to the model’s weights. Although this technique is promising, it necessitates considerable computational investment for hypernetwork training and frequently diminishes in effectiveness with the increase in language model size [30]. Optimization method ROME [19] employs causal mediation analysis to identify the editing region and focuses on altering specific information via rank-one adjustments to individual matrices. MEMIT [20] adhered to a similar methodology, adeptly modifying several parameter matrices concurrently to facilitate the simultaneous alteration of 10,000 knowledge entities, and showcasing robust generalization and specificity. PMET [15] advanced this technique, refining MEMIT’s capabilities for more precise editing. However, previous research indicates that minor modifications to the parameters of large language models can impact the model’s ultimate behaviour [25], and these methods do not allow the model to use new knowledge for reasonable inference. This means a user cannot selectively apply edits with these methods [13, 23].
@@
To solve the problems of existing knowledge editing methods and enhance the application of prompt engineering in the field of knowledge editing and inferencing, we proposed the PSPEM (Prefix Soft-Prompt Editing Method), an innovative strategy rooted in prompt engineering that can be used for a lifetime with just one training. This method allows for precise, nuanced modifications to the model’s output by employing a single knowledge prompt, all without altering the model’s parameters. Specifically, PSPEM utilizes a prompt encoder and an encoding converter to meticulously refine the prompt’s essential information. It ensures the accuracy of the model’s output by maximizing the probability assigned by the language model to the target token, and by aligning with the original prompt’s influence on continuation words, it guarantees the fluency of the output text and a high degree of consistency between the generated text information and the prompt information. We conducted evaluations on two knowledge editing tasks and two attribute inserting tasks. In the tasks of knowledge editing, PSPEM achieved nearly 100% editing accuracy while ensuring the fluency and consistency of the generated text. In terms of attribute inserting, PSPEM reached the state-of-the-art. The model can make reasonable inferences using the given prompts and generate text that aligns with the prompt information. We then conducted an analysis of the parallels between PSPEM and original prompts, measuring their impact on model output from multiple perspectives. The experimental results indicate that the impact of PSPEM on the model is highly similar to that of the original prompts, therefore, PSPEM can serve as an alternative to original prompts, supporting the model in effective editing. We summarize our contributions as follows:
@@
    We propose PSPEM, a lifetime knowledge editing method based on soft prompts that corrects the model’s output by learning information from the original prompts. In terms of editing effectiveness, it is superior to all other existing weight-reserved methods.
@@
    We evaluated PSPEM on two mainstream datasets for knowledge editing and two datasets for attribute inserting. The experimental results show that PSPEM can not only perform efficient and accurate knowledge editing but also utilize the given prompts for reasonable inference, which is beyond the capabilities of traditional prompt engineering and other knowledge editing methods.
@@
    As far as we know, PSPEM was the first attempt to use soft prompts for model knowledge editing and inference, which provides a feasible research direction for knowledge editing.
@@ 2.1 Preliminaries
This study centers on enhancing the application of prompt engineering in the field of knowledge editing and inferencing. As mentioned earlier, while prompt engineering enables models to adapt to diverse tasks without retraining, the search for suitable prompts is time-consuming and laborious. More importantly, when the information in the prompt conflicts with the internal knowledge of the model, such prompts often lose their effectiveness. This situation is demonstrated in Figure 1, where the model is prompted: “Banana has four legs” (left side of 1), and the model responds: “Banana is a fruit” (upper right side of 1). Clearly, the prompt has lost its effect, with the model still recognizing the banana as a fruit, not an animal. PSPEM addresses this issue by enabling the model to correctly respond: “Banana is an animal.” Note that PSPEM does not cause confusion within the model, but rather makes the model pay more attention to the information in the prompts.
@@
Here, $W^{l}_{\text{MHSA}}$ and $W_{proj}$ refer to the output weights of the MHSA and MLP at the $l$-th layer, respectively, while $\sigma$ denotes the non-linear activation function. Different LMs frequently exhibit slight variations in implementing these transformations. Our goal is not to provide a full survey of these details but to capture essential terminology for our results.
@@ 2.2 PSPEM
PSPEM focuses on extracting key information from the original prompt and makes the subsequent text pay more attention to the prompt information. The overview of our proposed method is shown in Figure 2. The PSPEM consists of Prompt Encoding, Encoding Converter, and Aligning Technology.
As shown in the bottom left of Figure 2, PSPEM starts with a given prompt, such as “Danielle Darrieux was born in America”, denoted as $p_{1:n}$. We obtain the embedded representation of the prompt through a Prompt Encoding mechanism. Some studies indicate that models based on the Transformer architecture can extract sentence representations [9, 14]. We focus on the GPT architecture, inputting the prompt $p_{1:n}$ into the origin model (GPT2-XL or GPT-J-6B), and take the output of a certain layer of the last token $h_{n}^{l}$ as compressed sentence representation.
The compressed sentence representation is processed through an Encoding Converter mechanism to obtain a more accurate sentence representation. These precise sentence representations can be considered a series of word embeddings, intended to make the model pay more attention to the key information in the prompt. The Encoding Converter is initiated through two Multi-Layer Perceptron (MLP) and an activation function GLUE. If we denote the dimension of $h$ as $d$, then the sizes of the two multilayer perceptrons are $W_{1}^{d\times d}$ and $W_{2}^{d\times d*3}$. This process is expressed with a formula as follows:
@@
We denote the continuation words as $c_{1:m}$, which in Figure 2 is “The mother tongue of Danielle Darrieux is”. We freeze the parameters of the original model and use alignment techniques to train the Encoding Converter mechanism.
@@ 3.1.1 Concepts
    Specificity: Editing should be implemented locally, and knowledge beyond the scope of editing should not be changed. For example, the prediction of "The president of Russia is" should be "Vladimir Putin", not "Emmanuel Macron".
@@ 3.1.2 Datasets
We chose ZsRE [7] and COUNTERFACT [19] as our foundational datasets. These two datasets are the most widely used in the field of knowledge editing, and almost all editing methods have been evaluated on these two datasets. To facilitate the comparison between different methods, we chose these. ZsRE is a question-answering dataset, each example contains a sentence that needs to be edited, paraphrase sentences generated by back-translation, and a sentence unrelated to the edited. The COUNTERFACT dataset is curated from Wikipedia and stands out as a rigorous benchmark tailored for GPT-like causal language models, presenting a challenging set of editing tasks, that allow us to distinguish superficial changes in wording from deeper changes that represent a meaningful change in fact. It contains over 21,000 records, each with different relations and entities, with the primary goal of editing knowledge by changing the object while keeping the subject and relation constant. The dataset includes not only paraphrase sentences, but also sentences unrelated to the knowledge to be edited in order to effectively discriminate between minor word changes, with particular emphasis on counterfactual scenarios.
@@ 3.1.3 Configurations
We split the ZsRE dataset into 70% for training, 10% for validation, and 20% for testing. For COUNTERFACT, we used 4500 instances for training, 500 for validation, and 5000 for testing. We use one paraphrase sentence from each instance as knowledge prompts and $h_{n}^{12}$ (GPT-J) and $h_{n}^{24}$ (GPT2-XL) are used to compress prompt representation, as with REMEDI [13]. Setting $\lambda_{1}$ = 1 and $\lambda_{2}$ = 1, with an initial learning rate of 1e-3, employing the Adam optimizer with a Linear Learning Rate Decay strategy, and stopped after the validation set accuracy did not improve in 3 epochs. All models were trained and reasoned on NVIDIA A100 40G GPUs.
@@ 3.1.4 Baselines
The baselines for comparison include direct weight-preserved methods:
@@
    PREFIX PROMPT uses a paraphrased sentence to guide the model in making knowledge modifications.
@@
    REMEDI [13] works by extracting attribute information from the prompt and then injecting it into the subject word via a linear transformation.
@@
    KE [7] develops an LSTM sequence model, which employs gradient information to predict the rank-1 weight alterations in the model. We resort to using the re-implemented version provided by Mitchell et al. [21] in their research.
@@
    MEND [21] is based on KE, adeptly manipulates the gradient of fine-tuned language models by capitalizing on a low-rank decomposition of the gradients, thereby enhancing the efficiency of the editing process.
@@
    MEMIT [20] builds upon ROME to insert many memories by modifying multi-layer MLP weights of a range of critical layers.
In fact, due to the rapid development in this research direction, some new editing methods have been proposed, like MALMEN [26] and PMET [16]. But these newly proposed methods are all based on MEND or ROME, lack some evaluation metrics and were not significant improvement in performance, so we do not include them in the baseline.
@@ 3.1.5 Metrics
    Neighborhood Score (NS) measures the $P_{LM}(o^{*})>P_{LM}(o^{c})$ of neighbourhood sentences that un-related to the knowledge that needs to be edited.
@@
Table 1 presents the performance of four weight-preserved editing methods. All these editing methods require some additional resources to assist model editing. When considering the NS metric, we set them all to 100, similar to Herandez et al [13].
PSPEM demonstrates the most outstanding performance among these four methods, with editing success rates approaching 100%. Compared to the prefix prompt method in 16 metrics, PSPEM only slightly underperforms on one RS metric in COUNTERFACRT, indicating its success in extracting critical information from prompts and effectively applying it to guide model outputs.
As an incremental step in prompt engineering, IKE achieved the highest ES metric in COUNTERFACT using the GPT-J model and surpassed PSPEM in RS. Although IKE attempts to improve model output through the use of multiple prefix prompts with promising results, it still falls short of PSPEM and it must provide multiple examples to guide each editing instance.
On the other hand, REMEDI, another attempt to extract key information from prefix prompts, captures only partial attribute information while ignoring the entire contextual information. PSPEM surpasses REMEDI in all aspects by extracting complete information from prefix prompts and aligning the refined information with the prefix prompts.
@@
In contrast, PSPEM still has strong competitiveness. Compared to KE, PSPEM significantly outperforms KE in metrics such as ES and GS. In tests using GPT2-XL on the ZsRE and COUNTERFACT datasets, PSPEM’s editing accuracies were 152% and 902% of KE’s, and the fluency of the generated text (GS) was significantly higher than KE’s.
Although PSPEM does not surpass editing success rates like MEND and ROME, it achieves the best performance in RS and GS. This indicates that PSPEM not only conducts reasonable editing but also effectively extends to the neighborhood and produces coherent and fluent text. Specific examples can be found in 3.3.

### v2

# Editing Knowledge Representation of Language Model via Rephrased Prefix Prompts

@@ Abstract
Neural language models (LMs) have been extensively trained on vast corpora to store factual knowledge about various aspects of the world described in texts. Current technologies typically employ knowledge editing methods or specific prompts to modify LM outputs. However, existing knowledge editing methods are costly and inefficient, struggling to produce appropriate text. Additionally, prompt engineering is opaque and requires significant effort to find suitable prompts. To address these issues, we introduce a new method called PSPEM (Prefix Soft-Prompt Editing Method), that can be used for a lifetime with just one training. It resolves the inefficiencies and generalizability issues in knowledge editing methods and overcomes the opacity of prompt engineering by automatically seeking optimal soft prompts. Specifically, PSPEM adopts a prompt encoder and an encoding converter to compress and refine key information in prompts and adopts prompt alignment techniques to guide model generation, ensuring text consistency and adherence to the intended structure and content. We have validated the effectiveness of PSPEM through knowledge editing and attribute inserting. On the COUNTERFACT dataset, PSPEM achieved nearly 100% editing accuracy and demonstrated the highest level of fluency. We further analyzed the similarities between PSPEM and original prompts and their impact on the model’s internals. The results indicate that PSPEM can serve as an alternative to original prompts, supporting the model in effective editing.
@@ 1 Introduction
Figure 1: By inputting the prompts on the left side into the model, traditional prompt engineering generate erroneous text, while PSPEM can correct such errors.
Language models based on the Transformer architecture, such as GPT and BERT, have revolutionized various natural language processing tasks with their capacity to store and utilize real-world factual knowledge within their parameters [27, 8, 24]. For example, when asked, “Where is the Eiffel Tower located?” GPT provides the accurate answer, “Paris”. However, inconsistencies or biases present in the pre-training data can propagate into the text generated by the model, leading to errors or contradictions. Additionally, the knowledge of the world changes over time, which presents a challenge to the static knowledge in the model. Addressing this issue requires a nuanced approach to updating the model’s knowledge base. Merely retraining the entire model with new data is prohibitively expensive and time-consuming, while fine-tuning, focusing solely on specific updated knowledge poses the risk of overfitting and compromising the model’s ability to generalize.
@@
Methods of model editing fall into two distinct classifications depending on the alteration of the original model weights: weight-preserved and weight-modified methods [30]. Weight-preserved strategies typically necessitate the inclusion of extra content, while weight-modified techniques directly alter the model’s weights. Weight modification methods include hypernetwork-based learning methods and direct optimization methods. Hypernetwork-based learning methods, such as KE [7], MEND [21] and MALMEN [26], utilize a hypernetwork to predict essential updates to the model’s weights. Although this technique is promising, it necessitates considerable computational investment for hypernetwork training and frequently diminishes in effectiveness with the increase in language model size [30]. Optimization method ROME [19] employs causal mediation analysis to identify the editing region and focuses on altering specific information via rank-one adjustments to individual matrices. MEMIT [20] adhered to a similar methodology, adeptly modifying several parameter matrices concurrently to facilitate the simultaneous alteration of 10,000 knowledge entities, and showcasing robust generalization and specificity. PMET [15] advanced this technique, refining MEMIT’s capabilities for more precise editing. However, previous research indicates that minor modifications to the parameters of large language models can impact the model’s ultimate behaviour [25], and these methods do not allow the model to use new knowledge for reasonable inference [13, 23].
@@
To overcome the limitation of the weight-modified method in utilizing new knowledge for reasoning and the poor editing accuracy and laborious of the weight-preserved method, we proposed PSPEM (Prefix Soft-Prompt Editing Method), an innovative strategy rooted in prompt engineering that can be used for a lifetime with just one training. This method allows for precise, nuanced modifications to the model’s output by employing a single knowledge prompt, all without altering the model’s parameters. Specifically, PSPEM utilizes a prompt encoder to extract information, an encoding converter to refine key information in prompts, and adopts prompt alignment techniques to guide model generation. It ensures the accuracy of the model’s output by maximizing the probability assigned by the language model to the target token, and by aligning with the original prompt’s influence on continuation words, it guarantees the fluency of the output text and a high degree of consistency between the generated text information and the prompt information. We conducted evaluations on two knowledge editing tasks and two attribute inserting tasks. In the tasks of knowledge editing, PSPEM achieved nearly 100% editing accuracy while ensuring the fluency and consistency of the generated text. In terms of attribute inserting, PSPEM reached the state-of-the-art. The model can make reasonable inferences using the given prompts and generate text that aligns with the prompt information. We then analyzed the parallels between PSPEM and original prompts, measuring their impact on model output from multiple perspectives. The experimental results indicate that the impact of PSPEM on the model is highly similar to that of the original prompts, therefore, PSPEM can serve as an alternative to original prompts, supporting the model in effective editing. We summarize our contributions as follows:
@@
    We propose PSPEM, a lifetime knowledge editing method based on soft prompts that corrects the model’s output by learning information from the original prompts.
@@
    We evaluated PSPEM on two mainstream datasets for knowledge editing and two datasets for attribute inserting. The experimental results show that PSPEM can not only perform efficient and accurate editing but also utilize the given prompts for reasonable reasoning, which is beyond the capabilities of traditional prompt engineering and other knowledge editing methods.
@@
    As far as we know, PSPEM was the first attempt to adopt soft prompts for model knowledge editing and inference, providing a feasible solution for the development of more intuitive and accurate language model editing tools.
@@ 2.1 Preliminaries
This study centers on enhancing the application of prompt engineering in the field of knowledge editing and inferencing. As mentioned earlier, while prompt engineering enables models to adapt to diverse tasks without retraining, the search for suitable prompts is time-consuming and laborious. More importantly, when the information in the prompt conflicts with the internal knowledge of the model, such prompts often lose their effectiveness. This situation is demonstrated in Figure 1, where the model is prompted: "Banana has four legs" (left side of 1), and the model responds: "Banana is a fruit" (upper right side of 1). The prompt has lost its effect, with the model still recognizing the banana as a fruit, not an animal. PSPEM addresses this issue by enabling the model to correctly respond: "Banana is an animal." Note that PSPEM does not cause confusion within the model, but rather makes the model pay more attention to the information in the prompts.
@@
Here, $W^{l}_{\text{MHSA}}$ and $W^{l}_{proj}$ refer to the output weights of the MHSA and MLP at the $l$-th layer, respectively, while $\sigma$ denotes the non-linear activation function. Different LMs frequently exhibit slight variations in implementing these transformations. Our goal is not to provide a full survey of these details but to capture essential terminology for our results.
@@ 2.2 PSPEM
PSPEM focuses on extracting key information from the original prompt and making the subsequent text and prompts more consistent. The overview of our proposed method is shown in Figure 2. The PSPEM consists of Prompt Encoding, Encoding Converter, and Aligning Technology.
As shown in the bottom left of Figure 2, PSPEM starts with a given prompt, such as "Danielle Darrieux was born in America", denoted as $p_{1:n}$. We obtain the embedded representation of the prompt through a Prompt Encoding mechanism. Some studies indicate that models based on the Transformer architecture can extract sentence representations [9, 14]. We focus on the GPT architecture, inputting the prompt $p_{1:n}$ into the origin model (GPT2-XL or GPT-J-6B), and take the output of a certain layer of the last token $h_{n}^{l}$ as compressed sentence representation.
The compressed sentence representation is processed through an Encoding Converter mechanism to obtain a more accurate sentence representation. These precise sentence representations can be considered a series of word embeddings, intended to make the model more focused on key information in the prompt. The Encoding Converter is initiated through two Multi-Layer Perceptron (MLP) and an activation function GLUE. If we denote the dimension of $h$ as $d$, then the sizes of the two multilayer perceptrons are $W_{1}^{d\times d}$ and $W_{2}^{d\times d*3}$. This process is expressed with a formula as follows:
@@
We denote the continuation words as $c_{1:m}$, which in Figure 2 is "The mother tongue of Danielle Darrieux is". We freeze the parameters of the original model and use alignment techniques to train the Encoding Converter.
Algorithm 1 Prefix Soft-Prompt Editing Method (PSPEM)
0:  Prompt sentence $p_{1:n}$, continuation words $c_{1:m}$, target token $t_{\text{target}}$.
0:  Edited text that aligns with the prompt.
1:  Initialize: Prompt Encoder, Encoding Converter;
2:  Input the prompt sentence $p_{1:n}$ into the model;
3:  Obtain last token representation $h_{n}^{l}$;
4:  Apply the Encoding Converter to transform $h_{n}^{l}$ into $h_{\text{emb}}$;
5:  Construct enhanced embedding $E_{s}$ by concatenating $h_{\text{emb}}$ and $c_{1:m}$;
6:  Input $E_{s}$ into the model to obtain outputs $[D_{h_{\text{e}mb}};D_{c_{1:m}}]$;
7:  Construct original embedding $E_{s}^{{}^{\prime}}$ by concatenating $p_{1:n}$ and $c_{1:m}$;
8:  Input $E_{s}^{{}^{\prime}}$ into the model to obtain outputs $[D^{{}^{\prime}}_{p_{1:n}};D^{{}^{\prime}}_{c_{1:m}}]$;
9:  Objective Function:        $L_{\text{target}}\leftarrow-P_{\mathrm{LM}}\left(c_{m+1}=t_{\mathrm{target}}\mid E_{s}{}\right)$;        $L_{\text{KL}}\leftarrow\sum_{1}^{m}KL\left(D_{c_{i}}\|D^{{}^{\prime}}_{c_{i}}\right)$;
10:  Train the Encoding Converter by optimizing $\mathcal{L}=\lambda_{1}L_{\text{target}}+\lambda_{2}L_{\text{KL}}$;
11:  Guide the model’s output using the Prompt Encoder and Encoding Converter.
12:  return  Edited text that aligns with the prompt.
@@
Once we have trained the Encoding Converter, we can utilize the Prompt Encoding and Encoding Converter to extract crucial information from the prompt, guiding the model’s output. See Algorithm 1 for the pseudo-code of PSPEM.
@@ 3.1.1 Concepts
    Specificity: Editing should be implemented locally, and knowledge beyond the scope of editing should not be changed. The prediction of "The president of Russia is" should be "Vladimir Putin", not "Emmanuel Macron".
@@ 3.1.2 Datasets
We chose ZsRE [7] and COUNTERFACT [19] as our foundational datasets. These two datasets are the most widely used in the field of knowledge editing, and almost all editing methods have been evaluated on these two datasets. To facilitate the comparison between different methods, we chose these. ZsRE is a question-answering dataset, each example contains a sentence that needs to be edited, paraphrase sentences generated by back-translation, and a sentence unrelated to the edited. The COUNTERFACT dataset is curated from Wikipedia and stands out as a rigorous benchmark tailored for GPT-like causal language models, presenting a challenging set of editing tasks, that allow us to distinguish superficial changes in wording from deeper changes that represent a meaningful change. It contains over 21,000 records, each with different relations and entities, with the primary goal of editing knowledge by changing the object while keeping the subject and relation constant. The dataset includes not only paraphrase sentences but also sentences unrelated to the knowledge to be edited to effectively discriminate between minor word changes, with particular emphasis on counterfactual scenarios.
@@ 3.1.3 Configurations
We split ZsRE into 70% for training, 10% for validation, and 20% for testing. For COUNTERFACT, we used 4500 instances for training, 500 for validation, and 5000 for testing. We use one paraphrase sentence from each instance as knowledge prompts and $h_{n}^{12}$ (GPT-J) and $h_{n}^{24}$ (GPT2-XL) are used to compress prompt representation, as with REMEDI [13]. Setting $\lambda_{1}$ = 1 and $\lambda_{2}$ = 1, with an initial learning rate of 1e-3, employing the Adam optimizer with a Linear Learning Rate Decay strategy, and stopped after the validation set accuracy did not improve in 3 epochs. All models were trained and reasoned on NVIDIA A100 40G GPUs.
@@ 3.1.4 Baselines
The methods for comparison include direct weight-preserved methods:
@@
    PREFIX PROMPT adopts a paraphrase sentence to guide the model in making knowledge modifications.
@@
    REMEDI [13] works by extracting attribute information from the prompt and then injecting it into the subject word via a linear transformation. Similar to Figure 1, REMEDI attempts to extract information from "born in America" rather than the entire sentence and injecting it into "Danielle Darrieux".
@@
    KE [7] develops an LSTM sequence model, which employs gradient information to predict the rank-one weight alterations in the model. We resort to using the re-implemented version provided by Mitchell et al. [21] in their research.
@@
    MEND [21] is based on KE, adeptly manipulates the gradient of fine-tuned language models by capitalizing on a low-rank decomposition of the gradients, thereby enhancing the accuracy of the editing process.
@@
    MEMIT [20] builds upon ROME to insert many memories by modifying MLP weights of a range of critical layers.
In light of the rapid advancements in editing methodologies, several novel approaches have emerged, including MALMEN [26] and PMET [16]. These methods extend upon foundational works such as MEND and ROME. However, upon thorough review, we found that they lack comprehensive evaluation across key metrics critical to our study’s aims, such as a lack evaluation in ZsRE. Furthermore, while these methods contribute to the field’s development, our preliminary analysis indicated that their performance improvements were not substantial enough to meet our criteria for a significant advancement. This decision was made to ensure a focused and rigorous evaluation within the scope of our research, though we acknowledge the potential of these methodologies in contributing valuable insights to the field.
@@ 3.1.5 Metrics
    Neighborhood Score (NS) measures the $P_{LM}(o^{*})>P_{LM}(o^{c})$ of neighborhood sentences that un-related to the knowledge that needs to be edited.
@@
Table 1 presents the performance of four weight-preserved editing methods. All these editing methods require some additional resources to assist model editing. When considering the NS metric, we set them all to 100, the same as Herandez et al [13].
PSPEM performs the best of these four methods, with editing success rates approaching 100%. Compared to the prefix prompt method on 16 metrics, PSPEM only slightly underperforms on one RS metric in COUNTERFACT, indicating its success in extracting critical information from prompts and effectively applying it to guide model output.
As an incremental step in prompt engineering, IKE achieved the highest ES metric in COUNTERFACT using the GPT-J model and surpassed PSPEM in RS. Despite IKE’s innovative approach of employing multiple prefix prompts, its practical application in editing is hampered by the challenge of pre-identifying suitable examples for model guidance. Conversely, PSPEM’s methodology, requiring only a single prompt for effective editing, offers a more feasible solution for lifelong editing endeavors.
On the other hand, REMEDI, another attempt to extract key information from prefix prompts, captures only partial attribute information while ignoring the entire contextual information. PSPEM surpasses REMEDI in all aspects by extracting complete information from prefix prompts and aligning the refined information with the prefix prompts. The subsequent ablation study, detailed in Experiment 3.4, will further elucidate the comparative of PSPEM and REMEDI.
@@
Notably, PSPEM demonstrated exceptional performance in comparison to KE (Knowledge Editing). On the ZsRE dataset, PSPEM achieved an editing success rate of 99.9%, vastly outperforming KE’s 65.6%. Similarly, on the COUNTERFACT dataset, PSPEM’s editing and prompt success rates reached 100%, significantly higher than KE’s 13.4% and 11.0%, respectively. Moreover, PSPEM excelled in generating text with superior fluency (GS) and consistency (RS), showcasing its comprehensive strength in both editing precision and output quality.
Furthermore, although PSPEM may not achieve editing success rates as high as MEND and ROME, it notably excels in RS. This suggests that PSPEM not only conducts proficient editing but also generates text that conforms to edited knowledge. Further details and specific examples can be found in Section 3.3.